# Beyond Two-Stage Training: Cooperative SFT and RL for LLM Reasoning

### v1

@@ 5 Related Work
6 Conclusion
Limitations
Although the BRIDGE framework effectively integrates the two-stage reasoning training process, several limitations remain. First, the reasoning traces in the SFT dataset are distilled from large-scale reasoning models, which may contain noise. Future work could leverage automated detection tools (Sun et al., 2025; Chen et al., 2023a) to identify and filter out such issues, potentially enhancing performance. Second, some studies (Zhou et al., 2024; Yu et al., 2025) have shown that the model’s reasoning capabilities can lack robustness. Incorporating robust fine-tuning techniques (Chen et al., 2023b, 2025b, 2025a) may help mitigate this limitation. Finally, the framework is evaluated in limited application scenarios. Future research could explore extending the BRIDGE framework to larger-scale models and broader domains, including program synthesis, theorem proving, and scientific reasoning.

### v2

@@ 5 Related Work
Recent progress has highlighted the critical role of reinforcement learning in enhancing the reasoning capabilities of large language models (OpenAI, ; DeepSeek-AI et al., 2025). Recent advances have been made in recipes for training reasoning models. SimpleRL (Zeng et al., 2025) observes that fine-tuning on short-CoT datasets can harm reasoning ability, while He et al. (2025) find that fine-tuning on long-CoT distilled data can improve the reasoning performance of smaller models—especially when used as a warm-up stage before RL training. In practice, two-stage pipelines that combine SFT and RL are commonly used to balance stability and performance. However, existing approaches often rely solely on supervised fine-tuning, which tends to generalize poorly, or on pure RL, which suffers from sample inefficiency and unstable optimization.
Recent efforts move beyond the decoupled “SFT then RL” recipe by mixing two objectives within one stage. Concurrent work CHORD (Zhang et al., 2025) integrates SFT into on-policy RL via a weighted-sum objective with a global weight and a token-level weight that up-weights uncertain expert tokens and down-weights large-divergence tokens. This stabilizes training compared with naïve loss addition, but the coupling is heuristic and offers no mechanism to ensure that the injected supervision is useful for the RL update. LUFFY (Yan et al., 2025) combines off-policy expert traces with on-policy rollouts using Mixed-Policy GRPO, but it requires the off-policy demonstrations to be paired with the same prompts as the on-policy data, limiting flexibility and data reuse. BRIDGE instead treats SFT–RL cooperation as a bilevel problem, meta-adapting the supervision to maximize the reward gain of joint training over RL alone and yields larger and more robust improvements than simple loss mixing. It offers a new perspective on integrating imitation and exploration for large reasoning models.
We provide an extended discussion of related work in Appendix E.
6 Conclusion
Appendix A The Use of Large Language Models
In preparing this manuscript, we used a large language model (LLM) solely for polishing the writing style and improving the clarity of the manuscript. The LLM was not used for generating research ideas, designing experiments, conducting analyses, or deriving results. All scientific contributions, including the conceptualization, methodology, experiments, and conclusions, were developed entirely by the authors.
Appendix B Implementation Details
[…]
Appendix C Performance on varied fine-tuning epochs
We assess BRIDGE’s effectiveness across different fine-tuning epochs on Qwen2.5-3B using average performance across epochs as the metric. As shown in Table 4.3, BRIDGE achieves the highest average performance. Among the baselines, Cold-start yields the second-best trade-off. However, its performance becomes unstable as training progresses, eventually converging to the same final result as RL-zero. In contrast, BRIDGE demonstrates consistent improvement throughout training. Overall, nearly all hybrid baselines outperform RL-zero in terms of early-stage efficiency, highlighting the advantage of integrating supervised fine-tuning and reinforcement learning paradigms.
Appendix D LoRA hyperparameter sensitivity
To ensure BRIDGE’s gains do not hinge on a particular LoRA setting and to assess robustness to LoRA. We ablate LoRA rank ($R$) and $\alpha$ on Qwen3-8B-Base, keeping all training settings fixed and only changing $(R,\alpha)$. Table 7 indicates nearly identical outcomes across configurations, confirming that BRIDGE is insensitive to the LoRA choice.
[…]
32/16 79.0 39.7 44.0 16.7 70.0 49.9
16/32 79.0 38.6 44.0 16.0 70.0 49.5
Table 7: LoRA sensitivity ablation on Qwen3-8B-Base.
Appendix E Related Work